OpenStudy (anonymous):

Two dice are rolled simultaneously. What is the probability that the product of the two dice is a multiple of 6?

OpenStudy (anonymous):

How do I solve for it step by step?

OpenStudy (jack1):

multiples of 6, between 6 and 36 are: 6, 12, 18, 24, 30, 36 so 6 possible answers

OpenStudy (anonymous):

list out all possible ways to get 6, 12, 18, 24, 30 or 36 by multiplying together the two numbers. the number of possible ways divided by the 36 (do you see why?) is the probability.

OpenStudy (anonymous):

So is the probability 5/12?

OpenStudy (anonymous):

@Peter14's idea is correct, but be sure to think of the two die as different color (say, red and green) because a R1 and a G6 is a different roll than a G1 and R6.

OpenStudy (jack1):

so for 1 die 6 and anything (6) 5 and 6 (1) 4 and 6 or 3 (2) 3 and any even number (3) 2 and 3 or 6 (2) 1 and 6 (1)

OpenStudy (anonymous):

here, i'll give you one example: all possible ways to get 6: g1*r6, g2*r3, g3*r2, g6*r1

zepdrix (zepdrix):

Jack I'm just a little worried that you're counting the same rolls twice with that first group of (6) and the ones below it.

OpenStudy (anonymous):

jack's way is the one you come up with when you can do it intuitively, my way is the procedural way

OpenStudy (jack1):

@zepdrix i think you're right

OpenStudy (anonymous):

I don't think jack's way is wrong... let's see.

OpenStudy (zarkon):

@danielle02 you are correct

OpenStudy (anonymous):

it is wrong...it double counts the 6/1 combination...i actually get 16/36 = 4/9. did i miss some combinations?

OpenStudy (anonymous):

but you should double count 6/1 because you can get 6/1 and 1/6

OpenStudy (anonymous):

the only one you don't double count is 6/6

OpenStudy (anonymous):

but he was only looking at a single die's values.

OpenStudy (zarkon):

(1,6) (2,6) (3,6) (4,6) (5,6) (6,6) (6,5) (6,4) (6,3) (6,2) (6,1) (2,3) (3,2) (2,4) (4,2)

OpenStudy (anonymous):

6/6 shouldn't be doubled because it is the same as itself

OpenStudy (zarkon):

(3,4) (4,3)

OpenStudy (anonymous):

@zrkon, why didn't you double count 6/6

OpenStudy (zarkon):

not(2,4)(4,2)

OpenStudy (zarkon):

there is only one way to get (6,6) you get a 6 and a 6 ;)

OpenStudy (anonymous):

but the dice are each independent. otherwise there'd be 35 possible combinations total

zepdrix (zepdrix):

Ahh i counted (6,6) twice :) No wonder I was getting 16 lol

OpenStudy (jack1):

so if the rest was right, 14=14 =28/36 = 7/9...?

OpenStudy (anonymous):

tht's why you think of them as different colors

OpenStudy (zarkon):

15/36=5/12

OpenStudy (anonymous):

so guys, what are we arguing about now?

OpenStudy (jack1):

@Zarkon that just seems a little low...?

OpenStudy (zarkon):

it is what it is

OpenStudy (anonymous):

jack, you don't multiply it by 2

OpenStudy (anonymous):

thanks guys

OpenStudy (jack1):

what about 3,4 and 4,3...?

OpenStudy (jack1):

shouldnt that make it 17/36...?

OpenStudy (anonymous):

if you want to do it your way, do it like this stem and leaf plot 6 | 123456 5 | 6 4 | 36 3 | 246 2 | 36 1 | 6 count the total number of leaves, and that's your number over 36

OpenStudy (anonymous):

apparently it's too early for me to do probability :P ignore my 6/6 discussion above :D

OpenStudy (anonymous):

you see I have covered every way to multiply to get a multiple of 6...

OpenStudy (jack1):

(1,6) (2,3) (2,4) (2,6) (3,6) (3,2) (3,4) (4,6) (4,3) (4,2) (5,6) (6,6) (6,5) (6,4) (6,3) (6,2) (6,1) ... this is the list of combo's... yeah?

OpenStudy (jack1):

and there are 17...?

OpenStudy (anonymous):

2,4 and 4,2 don't work...product is 8

OpenStudy (jack1):

ah ha , lol my bad